### v2

@@ 3 The Haemers bound and the minrank parameter
Let $\mathcal{F}\subset\binom{[n]}{k}$ and let $q$ be a prime power. Suppose that for all distinct $F,F^{\prime}\in\mathcal{F}$, $|\mathcal{F}\cap\mathcal{F}^{\prime}|\not\equiv k\pmod{q}$. Then,
@@ 4 Explicit constructions of large gaps between the Shannon capacity and the Lovász number
Given $\epsilon>0$, choosing $q$ sufficiently large implies the second part of Theorem 1.6. The first part of Theorem 1.6 is a consequence of the second part of the theorem and the inequality $\alpha(G_{q}(n,k))\leq c(G_{q}(n,k))$. The third part follows from the inequality $\chi(G_{q}(n,k))\geq\frac{N}{\alpha(G_{q}(n,k))}=\Omega(n^{q^{2}-q})$ and the fact that $\vartheta(G_{q}(n,k)^{C})=\Theta(n^{q-1})$, which follows from Theorem 2.5. (Well-known properties of the distribution of primes imply that the result for all sufficiently large integers $n$ follows from the construction presented here). ∎
@@
For any $\epsilon>0$ and any prime $p$, there is an explicit construction of a graph $G$ on $n$ vertices such that

### v3

@@ 3 The Haemers bound and the minrank parameter
Let $\mathcal{F}\subset\binom{[n]}{k}$ and let $q$ be a prime power. Suppose that for all distinct $F,F^{\prime}\in\mathcal{F}$, $|F\cap F^{\prime}|\not\equiv k\pmod{q}$. Then,
@@ 4 Explicit constructions of large gaps between the Shannon capacity and the Lovász number
Given $\epsilon>0$, choosing $q$ sufficiently large implies the second part of Theorem 1.6. The first part of Theorem 1.6 is a consequence of the second part of the theorem and the inequality $\alpha(G_{q}(n,k))\leq c(G_{q}(n,k))$. The third part follows from the inequality $\chi(G_{q}(n,k))\geq\frac{N}{\alpha(G_{q}(n,k))}=\Omega(n^{q^{2}-q})$ and the fact that $\vartheta(G_{q}(n,k)^{C})=\Theta(n^{q-1})$, which follows from Theorem 2.5. ∎
@@
For any $\epsilon>0$ and any prime $p$, for infinitely many $n$ there is an explicit construction of a graph $G$ on $n$ vertices such that
@@
We finally note that Theorem 1.6 also holds if the Lovász number is replaced by the Schrijver variant $\sigma$ (or the Delsarte linear programming bound). In particular, the graphs $G_{q}(n,k)$ again give an explicit construction of a graph $G$ on $N$ vertices with $\sigma(G)/c(G)=\Omega(N^{1-\epsilon})$.